1. Evaluate the following sums given the sequence: 16, 14, 30, 44,...
F(1) + F(2) + F(3) = ______
F(1) + F(2) + F(3) + F(4) = _____
F(1) + F(2) + F(3) + F(4) + F(5) = _____
2. Determine the pattern in the successive sums from the previous question. What will be the sum of F(1) + F(2) + … + F(10)?
3. If you have a wooden board that is 0.75 meters wide, how long should you cut it such that the Golden Ratio is observed? Use 1.618034 as the value of the Golden Ratio. (

Answer by ikleyn(52787)

.

H I N T


Notice that the terms starting from the third, are the sum of the two preceding terms.


It gives you the recursive formula   F(n+1) = F(n-1) + F(n),  n = 2, 3, 4, . . . 



Having this HINT, you can complete the assignment on your own.